# PromptDresser: Improving the Quality and Controllability of Virtual Try-On via Generative Textual Prompt and Prompt-aware Mask

### v1

@@ Abstract
Recent virtual try-on approaches have advanced by fine-tuning the pre-trained text-to-image diffusion models to leverage their powerful generative ability. However, the use of text prompts in virtual try-on is still underexplored. This paper tackles a text-editable virtual try-on task that changes the clothing item based on the provided clothing image while editing the wearing style (e.g., tucking style, fit) according to the text descriptions. In the text-editable virtual try-on, three key aspects exist: (i) designing rich text descriptions for paired person-clothing data to train the model, (ii) addressing the conflicts where textual information of the existing person’s clothing interferes the generation of the new clothing, and (iii) adaptively adjust the inpainting mask aligned with the text descriptions, ensuring proper editing areas while preserving the original person’s appearance irrelevant to the new clothing. To address these aspects, we propose PromptDresser, a text-editable virtual try-on model that leverages large multimodal model (LMM) assistance to enable high-quality and versatile manipulation based on generative text prompts. Our approach utilizes LMMs via in-context learning to generate detailed text descriptions for person and clothing images independently, including pose details and editing attributes using minimal human cost. Moreover, to ensure the editing areas, we adjust the inpainting mask depending on the text prompts adaptively. We found that our approach, utilizing detailed text prompts, not only enhances text editability but also effectively conveys clothing details that are difficult to capture through images alone, thereby enhancing image quality. Extensive experiments demonstrate that our method outperforms the baselines significantly while showing versatile manipulation capabilities based on text prompts. Our code is available at https://github.com/rlawjdghek/PromptDresser.
Figure 1: Generated results of PromptDresser: PromptDresser (a) enables high-quality virtual try-on through rich text description, (b) highlights multiple attributes manipulation simultaneously, and (c) generates versatile outputs across multiple clothing categories, including bottoms and dresses.
@@ 1 Introduction
Virtual try-on [21] provides an advanced technique for personalized fashion previews, removing the need for physical trials. Moreover, the capability to control wearing styles, such as tucking style or fit, can enhance overall user experience by allowing users to visualize different outfit options.
@@
In this paper, we tackle a text-editable virtual try-on task that changes the clothing item based on the provided clothing image and edits the wearing style according to the text descriptions. Text-editable virtual try-on has unique and challenging aspects. (i) It is crucial to construct rich and well-aligned text descriptions for wearing the new clothing to the person image, (ii) avoiding textual conflicts between the original and new clothing during sampling. Since the wearing styles (e.g., untucked, fully tucked in) vary according to the the text descriptions, (iii) an adaptive mask aligned with the text prompt is necessary to preserve regions irrelevant to the clothing and to minimize the influence of the original clothing shape.
@@
To address these challenges, we introduce PromptDresser, a novel virtual try-on model that leverages rich textual information to achieve high-quality and versatile manipulation. To obtain the text prompts for paired person-clothing data, we instruct the large multimodal models (LMMs) to describe the person and clothing images, respectively. However, LMMs often produce excessively diverse text descriptions on the provided images, leading to potential misalignment with the images or insufficient details. Therefore, we leverage in-context learning that conditions expressive layouts and clothing, effectively specifying attributes to focus on the inpainting regions with minimal human cost. This approach significantly outperforms the model that relies on holistic descriptions of the entire person and clothing image, providing a scalable and efficient solution for versatile virtual try-on. Furthermore, we found that detailed text descriptions produced by LMMs not only enhance text editability but also effectively maintain clothing details that are difficult to capture through clothing images alone, thereby generating high-fidelity images.
Following the previous approaches [11, 29, 38, 28, 12, 10], we also utilize an inpainting mask to determine the preserved and generated regions. The traditional inpainting mask often constrains the model to the shape of the existing clothing, leading to unnatural results when adapting to different clothing types. To mitigate such constraints, we applied random dilation mask augmentation, enabling the model to learn from a wide range of mask sizes, from broad to narrow. Expanding the masked area effectively removes details related to the original clothing’s length and shape, but it also poses the risk of erasing areas that should be preserved (e.g., pants when generating a top). Therefore, we propose to use an expanded inpainting mask to obtain an approximate clothing region aligned with the text prompt. Next, we generate a refined mask by combining this expanded mask with a fine mask that removes areas where clothing will be applied (e.g., arms or torso) to preserve as much of the original person and background as possible. By employing a refined mask that is agnostic to the existing clothing and aligned with the text prompts, our model preserves details irrelevant to the worn clothing, enabling a wide range of text-based manipulation. Through extensive experiments, we demonstrate that PromptDresser achieves superior image quality compared to the existing virtual try-on methods while effectively controlling the wearing styles using diverse text prompts.
@@
  • We propose a text-editable virtual try-on model that, with the assistance of a large multimodal model (LMM) with in-context learning, achieves rich, well-aligned text descriptions for both person and clothing, ensuring no textual conflicts with any clothing provided.
  • To mitigate the issue of following the original clothing’s attributes, such as shape and length, we propose random dilation mask augmentation. Our prompt-aware mask generation enhances diversity in virtual try-on results while effectively preserving the person’s original appearance.
@@ 2.1 Image-based Virtual Try-On
Early approaches often relied on two-stage frameworks, combining explicit warping networks with GAN-based generators [52, 11, 17, 29, 3, 50]. However, these methods continue to struggle with error accumulation across multiple stages, leading to a shift towards diffusion models which have shown impressive generative capabilities across various domains [55, 43, 54, 6, 58]. Due to the challenges of dataset acquisition and the advantages of leveraging prior knowledge, most approaches build upon large-scale text-to-image diffusion models [44, 42, 39, 9], inherently benefiting from their robust inpainting capabilities [19] or utilizing textual inversion techniques [38]. Notably, recent research [28, 12] propose attention-based, end-to-end virtual try-on models that preserve fine details of clothing while achieving the generalization performance.
Another line of progress in image-based virtual try-on is controllability. LC-VTON [53], for instance, introduced new segmentation labels to incorporate clothing length, enabling the generation of high-fidelity images. Additionally, some studies [31, 10] have utilized landmarks to introduce point-based controllability. While these spatial condition-based approaches offer a fine degree of controllability, they are limited in their ability to perform comprehensive editing such as adjusting fit and overall appearance. On the other hand, recent research leveraging text-based controllability has tried to address such challenges. However, most existing studies still incorporate text but fail to consider the existing areas (e.g., background) that need to be preserved [59] or are limited to captioning solely for clothing [34, 12].
In this paper, we propose a virtual try-on model that uses rich text descriptions to harness the capabilities of large-scale text-to-image diffusion models. By independently extracting captions for both the person and the clothing with LMMs, we enhance generalization performance and enable manipulation. Moreover, we propose a novel adaptive mask for further flexible manipulation while preserving the original person’s appearance.
@@ 3.4 Enhancing Adaptability via Mask Refinement
Prompt-aware Mask Generation. For the inference stage, we introduce a novel coarse-to-fine generation approach to effectively preserve the original person’s appearance while allowing flexible text-based image manipulation. As illustrated on the left side of Fig. 3, we begin inference with a coarse mask, aiming to create an initial approximation of the clothing region aligned with the text prompt. For efficiency, we apply early stopping in the denoising process, running it only from timestep $T$ to $\sigma T$, where $\sigma\in[0,1)$. We then approximate $\hat{\mathbf{z}}_{0}$, decode it to $\hat{\mathbf{x}}_{0}$, and segment the region of interest using an off-the-shelf human parsing model. By taking the union of this mask with an existing clothing-agnostic person representation (fine mask), we acquire a refined mask, which subsequently serves as the inpainting mask for generating the final output. This method achieves a balance between precision and efficiency, improving the alignment of the output with the text prompt.
4 Experiment
Benchmarks. We train PromptDresser separately on VITON-HD [11] and DressCode [37] datasets, and SHHQ-1.0 is used to evaluate the generalizability of the model trained on VITON-HD [28].
We compare our model to two GAN-based models (HR-VITON [29], GP-VTON [50]) and four diffusion-based models (LADI-VTON[38], DCI-VTON [19], StableVITON [28], and IDM-VTON [12]). We use pre-trained weights if available; otherwise, we re-implement them using official code. LADI-VTON, DCI-VTON, and StableVITON, all based on Stable Diffusion 1.5, generate images at 512$\times$384 resolution. To ensure a fair comparison, we upscale the outputs to 2$\times$ using Real-ESRGAN [49].
@@ 4.1 Comparison with Baselines
Qualitative Results Fig. 4 shows a comparative analysis of our model and baseline models trained on the VITON-HD [11] dataset (first row), evaluated both on the VITON-HD test dataset and the SHHQ-1.0 [16] dataset (second row). The baselines either closely follow the original clothing shape or lack detailed clothing features (especially on SHHQ-1.0). In contrast, our model accurately captures the details and shape of the reference clothing while preserving the features of the original person image, such as background and pose. Additionally, we compared images of the upper body, lower body, and dresses on the DressCode [37] dataset in Fig. 5. Across these images, the baselines tend to follow to the original clothing shape. Specifically in the lower body, baselines generate unnatural jeans by following the existing pants shape even when provided with a short skirt; similarly, for dresses, models including OOTDiffusion and IDM-VTON produce excessively long dresses. This demonstrates that well-aligned and rich textual information enhances model generalization performance, and our adaptive mask disregard the original clothing shape but retain the person’s appearance.
@@
Ours w/ holi. 0.8440 0.1345 9.41 1.33
Ours w/o PMG 0.8262 0.1538 9.11 0.98
Ours 0.8530 0.116 8.64 0.75
@@
Quantitative Results For quantitative evaluation, we use four metrics: SSIM [13], LPIPS [56] for the paired setting, and FID [24], and KID [5] for the unpaired setting. Tables 1 and 2 show the comparison results between PromptDresser and the baseline models on the VITON-HD and DressCode datasets, respectively. Our model consistently and significantly outperforms existing models in the unpaired setting (i.e., FID and KID). Notably, applying the generative textual prompt to IDM-VTON, we observed performance improvements across all metrics except LPIPS, highlighting our method’s effectiveness and the benefit of leveraging textual information about person and clothing attributes to enhance the realism of the generated images.
@@
While we include pose descriptions and hand as hints, there is a slight performance drop in the paired setting due to minor misalignment in the pose. We also experimented with expanding the first convolution layer of our model to 13 channels to include DensePose as an additional input [28], named as Ourspose. Ourspose achieved the highest LPIPS score, as shown in Table 1. However, a slight decrease in performance is observed in the unpaired setting due to modifications in the backbone architecture, which, along with pose-related errors, indicates a drop in generalization performance. As shown in qualitative evaluations, visual differences were minimal, so we selected the model without spatial pose conditioning as our final choice, given its superior scores in the unpaired setting.
4.2 Further Analyis on PromptDresser
@@
Evaluation on Text Alignment To validate our method’s editing capability, we generated edited versions of 2,032 test images from VITON-HD by fixing a specific attribute (e.g., “tucking style”) to a caption (e.g., “untucked”) and then evaluated whether the captions generated by the LMM for these edited images matched the intended caption. We conducted experiments with two settings: (i) setting the tucking style to “untucked” (i.e., the top is worn outside the pants) and (ii) setting the clothing fit to “tight fit”. The “Base Ratio” is the proportion of the 2,032 test images in which the LMM (i.e., GPT-4o) identified a specific caption. For example, if the LMM identifies 1,016 out of the 2,032 test images to have an “untucked” tucking style, the base ratio for ‘untucked’ would be 50%. In addition to two baselines, LADI-VTON and IDM-VTON, we compared our method with Ourspose to examine the effect of DensePose information on text-editability with the same text prompt.
@@
Table 3 shows that our model achieved significantly higher accuracy for both attributes. Notably, the baseline models scored 50.78% and 46.31% for the “untucked”, performing similarly to the base ratio. This result underscores the limitations of conventional agnostic masks that restrict manipulation to pre-defined areas, especially for adjusting clothing length. Furthermore, for the “tight fit” attribute, IDM-VTON and ourspose both achieved around 44%, while our method reached 66.98%, indicating that DensePose can hinder accurate text-based editing. Therefore, the results demonstrate that our approach addresses the limitations of conventional agnostic mask, enabling for accurate manipulation based on text prompts.
@@
Ablation Study We investigate the key design choices of our method through ablation studies on the VITON-HD dataset, as shown in Fig. 6 and Table 4. Our comparisons include: (i) using a holistic text description (i.e., a overall description of the image in detail) for both person and clothing images with the LMM and (ii) excluding the prompt-aware mask generation (PMG) during inference.
Fig. 6 shows that using a holistic description can fail to capture complex poses accurately. Moreover, our method with holistic description shows the lowest FID and KID scores in Table 4. These results indicate that text prompts describing the images using LMM without in-context learning are insufficient to represent fine-grained details including poses. Since Ours without PMG employs only the coarse masks, the regions that need to be preserved are often generated differently, as shown in Fig. 6 (e.g., a belt absent on the target person). Table 4 demonstrate that PMG improves the performance significantly in all metrics.
@@
[…]
Limitations and Future Work This paper demonstrated the utility of generative text prompts using LMMs with in-context learning. However, due to the limitation of SDXL that accepts only up to 77 text tokens, we manually constructed in-context dataset. Inspired by recent studies [32] on configuring in-context sequences to enhance in-context learning performance, exploring in-context data construction for virtual try-on can be a promising research direction. Stable Diffusion 3 [15] has employed a T5 text encoder [40] capable of covering long context lengths. Utilizing well-configured in-context learning on such models is expected to further improve the performance of virtual try-on.

### v2

@@ Abstract
Recent virtual try-on approaches have advanced by fine-tuning pre-trained text-to-image diffusion models to leverage their powerful generative ability; however, the use of text prompts in virtual try-on remains underexplored. This paper tackles a text-editable virtual try-on task that modifies the clothing based on the provided clothing image while editing the wearing style (e.g., tucking style, fit) according to the text descriptions. In the text-editable virtual try-on, three key aspects exist: (i) designing rich text descriptions for paired person-clothing data to train the model, (ii) addressing the conflicts where textual information of the existing person’s clothing interferes the generation of the new clothing, and (iii) adaptively adjust the inpainting mask aligned with the text descriptions, ensuring proper editing areas while preserving the original person’s appearance irrelevant to the new clothing. To address these aspects, we propose PromptDresser, a text-editable virtual try-on model that leverages large multimodal model (LMM) assistance to enable high-quality and versatile manipulation based on generative text prompts. Our approach utilizes LMMs via in-context learning to generate detailed text descriptions for person and clothing images independently, including pose details and editing attributes using minimal human cost. Moreover, to ensure the editing areas, we adjust the inpainting mask depending on the text prompts adaptively. Our approach enhances text editability while effectively conveying clothing details that are difficult to capture through images alone, leading to improved image quality. Experiments show that PromptDresser significantly outperforms baselines, demonstrating superior text-driven control and versatile clothing manipulation. Our code is available at https://github.com/rlawjdghek/PromptDresser.
Figure 1: Generated results of PromptDresser: (a) effectively resolves the issue of residual clothing shape in existing methods (b) highlights multiple attributes manipulation simultaneously, and (c) generates versatile outputs across multiple clothing categories, including bottoms and dresses.
@@ 1 Introduction
Inpainting-based Virtual Try-on [21] effectively preserves non-target clothing features (e.g., face or background), enabling personalized fashion previews by swapping items such as tops and bottoms. Moreover, the capability to control wearing styles, such as tucking style or fit, can enhance overall user experience by allowing users to visualize different outfit options.
@@
In this paper, we tackle a text-editable virtual try-on task that changes the clothing item based on the provided clothing image and edits the wearing style according to the text descriptions. Text-editable virtual try-on has unique and challenging aspects. (i) Constructing rich, yet compact and well-aligned text descriptions for dressing the person in new clothing is essential, (ii) while avoiding textual conflicts between the original and new clothing during sampling. Since wearing styles (e.g., untucked, fully tucked in) should be adjustable through text, (iii) the mask must be adaptively change to align with the text prompt, preserving regions unrelated to clothing (e.g., face or background) and removing the information of the original clothing shape.
@@
To address these challenges, we introduce PromptDresser, a novel virtual try-on model that leverages rich textual information to achieve high-quality and versatile outfit manipulation. To generate the text prompts for paired person-clothing data, we instruct the large multimodal models (LMMs) to describe the person and clothing images, respectively. In virtual try-on, the target clothing differs from the originally worn clothing during inference. For a high-fidelity try-on, it is crucial to retain only agnostic information while removing details of the original clothing [10]. Additionally, unmasked regions in the inpainting process already contain significant details, such as the face and non-target clothing. Simply instructing an LMM to describe the image often produces mixed captions of person and clothing, or generate excessively redundant descriptions. Therefore, we leverage in-context learning that conditions expressive layouts, effectively specifying attributes to focus on the inpainting regions with minimal human cost. We pre-define attributes for inpainting-based virtual try-on and provide human-labeled attribute examples for LMM in-context learning. This approach is significantly effective than holistic descriptions of the entire person and clothing, providing a scalable and efficient solution for virtual try-on. Moreover, we find that detailed text descriptions generated by LMMs not only improve text-based editability but also preserve intricate clothing details that images alone may not fully capture, resulting in high-fidelity virtual try-on images.
In virtual try-on [10, 29, 38, 28, 11, 9], inpainting masks play a crucial role in removing existing clothing details. Notably, conventional masks often constrain the model to the original clothing shape, leading to unnatural results with different clothing types (see Figure 1 (a)). To address this, we introduce a novel masking technique, Prompt-aware Mask Generation (PMG), which adaptively adjusts according to the text prompt. We propose two types of mask: coarse and fine mask. The course mask is used to predict the form of the new clothing, agnostic to the existing clothing. It serves as a guide to indicate the areas needed for the new clothing to naturally generate while aligned with the text descriptions (e.g., outfit or tucking style). By combining the predicted mask and the fine mask, we adaptively mask out the existing clothing conditioned on the text prompt while preserve the non-target clothing region. To enable the model to learn from coarse to fine masks, we additionally applied random dilation mask augmentation during the training stage. This enables the model to learn how to adaptively generate across a wide range of mask regions, from broad to narrow, thereby improving its flexibility in handling various mask shapes.
@@
  • We propose a text-editable virtual try-on model that, with the assistance of a LMM with in-context learning, achieves rich, well-aligned text descriptions for both person and clothing, ensuring no textual conflicts with any clothing provided.
  • To mitigate the issue of following the original clothing’s attributes, such as shape, we propose random dilation mask augmentation. Our prompt-aware mask generation enhances diversity in virtual try-on results while effectively preserving the person’s original appearance.
@@ 2.1 Image-based Virtual Try-On
Recent virtual try-on approaches build upon large-scale text-to-image diffusion models [44, 43, 39, 8], inherently benefiting from their robust inpainting capabilities [18] or utilizing textual inversion techniques [38] or conditioning via attention mechanism [28, 11]. Notably, recent research [28, 11] propose attention-based, end-to-end virtual try-on models that preserve fine details of clothing while achieving the generalization performance.
Another line of progress in image-based virtual try-on is controllability. LC-VTON [53], for instance, introduced new segmentation labels to incorporate clothing length, enabling the generation of high-fidelity images. Moreover, while spatial condition-based approaches [31, 9] offer a fine degree of controllability, they are limited in their ability to perform comprehensive editing such as adjusting fit and overall appearance.
On the other hand, recent research leveraging text-based controllability has tried to address such challenges, but fail to consider the existing areas (e.g., background) that need to be preserved [57] or are limited to captioning solely for clothing [34, 11]. While some methods condition on both text and clothing [45], effective prompting and masking techniques for inpainting-based virtual try-on remain unexplored. In this paper, we propose an effective prompting strategy for disentangling the person and clothing in inpainting-based virtual try-on, along with a novel adaptive masking method that enables more flexible manipulation while preserving the original person’s appearance.
@@ 3.4 Enhancing Adaptability via Mask Refinement
Prompt-aware Mask Generation. For inference, we introduce a novel coarse-to-fine generation to effectively preserve the original person’s appearance while allowing flexible text-based image manipulation. As shown in Fig. 3, we start with a coarse mask to create an initial approximation of the clothing region aligned with the text prompt. For efficiency, we apply early stopping in the denoising process, running it only from timestep $T$ to $\sigma T$, where $\sigma\in[0,1)$. We then approximate $\hat{\mathbf{z}}_{0}$, decode it to $\hat{\mathbf{x}}_{0}$, and segment the region of interest using an off-the-shelf human parsing model. By taking the union of this mask with an existing clothing-agnostic person representation (fine mask), we acquire a refined mask, which subsequently serves as the inpainting mask for generating the final output. This approach balances precision and efficiency, improving the alignment with the text prompt.
@@
4 Experiment
We train PromptDresser separately on VITON-HD [10] and DressCode [37] datasets, and SHHQ-1.0 is used to evaluate the generalizability of the model trained on VITON-HD [28]. We provide details on the baseline and implementation in the supplementary material.
@@ 4.1 Comparison with Baselines
Qualitative Results. Fig. 4 compares images generated by a model trained on VITON-HD, evaluating its performance on both the VITON-HD dataset (first row) and the SHHQ-1.0 dataset (second row). The baselines either closely follow the original clothing shape or lack detailed clothing features (especially on SHHQ-1.0). In contrast, our model accurately captures the details and shape of the reference clothing while preserving the features of the original person image, such as background and pose. Additionally, we compared images of the upper body, lower body, and dresses on the DressCode [37] dataset in Fig. 5. Across these images, the baselines tend to follow to the original clothing shape. Specifically in the lower body, baselines generate unnatural jeans by following the existing pants shape even when provided with a short skirt; for dresses, baselines also produce excessively short dresses.
@@
Ours w/ holi. 0.8574 0.1296 9.21 1.18
Ours w/o PMG 0.8411 0.1485 8.98 0.87
Ours 0.8686 0.1119 8.54 0.67
@@
Quantitative Results. For quantitative evaluation, we use four metrics: SSIM [12], LPIPS [54] (paired setting), and FID [24], KID [5] (unpaired setting). Tables 1 and 2 show the comparison between PromptDresser and the baseline models on the VITON-HD and DressCode datasets, respectively. Our model consistently outperforms existing models in the unpaired setting (i.e., FID and KID). Notably, applying the generative textual prompt to IDM-VTON improved performance across all metrics except LPIPS, demonstrating the effectiveness of leveraging textual information for more realistic image generation.
@@
While adding pose and hand descriptions slightly impacted performance in the paired setting due to minor pose misalignment, expanding the first convolution layer to 13 channels incorporating DensePose [28] (Ourspose) shows the highest SSIM and LPIPS, as shown in Table 1. However, performance dropped in the unpaired setting due to backbone modifications and pose-related errors, affecting generalization. Given minimal visual differences in qualitative and better unpaired results, we selected the model without spatial pose conditioning as the final choice.
4.2 Further Analysis on PromptDresser
@@
Evaluation on Text Alignment. To validate our method’s editing capability, we generated edited versions of 2,032 test images from VITON-HD by fixing a specific attribute (e.g., “tucking style”) to a caption (e.g., “untucked”) and then evaluated whether the captions generated by the LMM for these edited images matched the intended caption. We tested two settings: (i) setting the tucking style to “untucked” (i.e., top worn outside the pants) and (ii) “tight fit”. The “Base Ratio” is the proportion of the 2,032 test images where the LMM identifies the caption. For example, if 1,016 images are identified as “untucked,” the base ratio for “untucked” is 50%. We compared our method with LADI-VTON, IDM-VTON, and Ourspose, all using text prompt.
@@
Table 3 shows that our model achieved significantly higher accuracy for both attributes. Notably, the baseline models scored 50.78% and 46.31% for the “untucked”, performing similarly to the base ratio. This result underscores the limitations of conventional agnostic masks that restrict manipulation to pre-defined areas, especially for adjusting clothing length. Furthermore, for the “tight fit” attribute, IDM-VTON and ourspose both achieved around 44%, while our method reached 66.98%, indicating that DensePose can hinder accurate text-based editing. These results show our approach outperforms conventional agnostic masks, enabling more precise manipulation from text prompts.
@@
Ablation Study. We investigate key design choices through ablation studies on the VITON-HD dataset, shown in Fig.6 and Table4. We compare: (i) using a holistic text description (i.e., a overall description of the image in detail) for both person and clothing images with the LMM, and (ii) excluding PMG during inference.
Fig.6 shows that a holistic description fails to capture complex poses accurately, and Table4 reveals it results in the worst FID and KID scores. This indicates that LMM-based prompts without in-context learning miss fine details like poses. Without PMG, the coarse masks often lead to inaccuracies (e.g., missing a belt, as shown in Fig.6). Table4 confirms that PMG significantly improves performance across all metrics.
Extended Applications of PromptDresser. In Figure 7, we further demonstrate our method’s effectiveness in layering outerwear. For instance, in the second row of of Fig. 7, both models were given a text: “woman is wearing a outerwear with the zipper open.”, and we use Grounded-SAM [42].
IDM-VTON struggle to accurately interpret the text prompt and rely on constrained mask regions during training and inference, causing them to conform to the shape of the existing innerwear. In contrast, our approach, enhanced by rich text prompts and PMG, successfully preserves the red innerwear, highlighting PMG’s ability to achieve precise, text-driven clothing manipulation.
Figure 7: Extended Applications of PromptDresser
[…]
Limitations and Future Work This paper demonstrated the utility of generative text prompts using LMMs with in-context learning. However, due to SDXL’s 77-token limit, we manually constructed the in-context dataset. Inspired by recent studies [32] on optimizing in-context sequences, exploring automated in-context data construction for virtual try-on is a promising direction. Stable Diffusion 3[14], with its T5 text encoder[40], supports longer contexts, and leveraging well-configured in-context learning on such models could further enhance virtual try-on performance.